OpenStudy (anonymous):

Please help 3(12 - 5) ^ 2 ÷ (4 + 3) = ?? Step by step

OpenStudy (anonymous):

Order of Operations: (PEMDAS) Parentheses Exponents Multiplication and Division (left to right) Addition and Subtraction (left to right)

OpenStudy (anonymous):

According to the order of operations: 1. Simplify what is inside the parentheses first: Can you do that and tell me what you got?

OpenStudy (anonymous):

3(12-5)^2 / (4+3) = ? 1. Simplify what is inside the parentheses first: 3(7)^2 / 7 Did you do this?

OpenStudy (anonymous):

Yes I did but keep getting different answers I just don't know what I'm doing wrong none of the answers I get are listed.

OpenStudy (anonymous):

I got 21 as the answer. Are you sure you wrote out the question correctly?

OpenStudy (anonymous):

I'm sorry I mean't the answer is 21 not 6 my bad.

OpenStudy (anonymous):

But how did you get that what am I doing wrong ?

OpenStudy (anonymous):

Ok let's go thru it step by step.

OpenStudy (anonymous):

THANKS SO MUCH :)

OpenStudy (anonymous):

please proceed

OpenStudy (anonymous):

1. Simplify the numbers in the parentheses. (12-5) = 7 (4+3) = 7. That leaves us with: 3(7)^2 / (7)

OpenStudy (anonymous):

Do u understand how i did that?

OpenStudy (anonymous):

is it 21?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

2. The next step is do the exponents. we only have one: 7^2 = 49. Now we have: 3*49/7

OpenStudy (anonymous):

okay so you multiply 49 times 3 and its 147 divide by 7 and hence 21 ?

OpenStudy (anonymous):

That's correct. Multiplication and Division goes from left to right

OpenStudy (anonymous):

Oh my god thank you so much I was really having suck a hard thank's a lot for your help and time ! :)

OpenStudy (anonymous):

such *

OpenStudy (anonymous):

Sure no problem. Do u have any other questions?

OpenStudy (anonymous):

Ummm I'm trying to figure them out on my own for now but once they get hard I'll be a bother and ask :P thanks a lot again !